You have just received a windfall from an investment you made in a​ friend's business. She will be paying you

$ 28,753 at the end of this​ year, $57,506 at the end of next​ year, and $86,259 at the end of the year after that​ (three years from​ today). The interest rate is 13.6% per year.

a. What is the present value of your​ windfall?

b. What is the future value of your windfall in three years​ (on the date of the last​ payment)?
Answer #1

a)

Present value = 28,753 / (1 + 0.136)1 + 57,506 / (1 + 0.136)2 + 86,259 / (1 + 0.136)3

Present value = 25,310.73944 + 44,561.16098 + 58,839.56115

Present value = $128,711.46

b)

Future value = 28,753(1 + 0.136)2 + 57,506(1 + 0.136)1 + 86,259(1 + 0.136)0

Future value = 37,105.63149 + 65,326.816 + 86,259

Future value = $188,691.45